# Gravitational four-derivative corrections in non-relativistic heterotic supergravity and the Green–Schwarz mechanism Eric Lescano

### v1

@@ Abstract
We present the first explicit construction of the four-derivative gravitational corrections to heterotic supergravity in the non-relativistic (NR) limit. By extending the Bergshoeff–de Roo identification to NR backgrounds, we obtain a finite higher-derivative completion of the NR heterotic action and determine its full set of symmetry transformations. A key outcome is the emergence of a gravitational Green–Schwarz mechanism from the NR gauge transformation of the Kalb–Ramond field. This mechanism can be trivialized through suitable field redefinitions, in agreement with a previous analysis of this transformation. Our results establish a systematic framework for incorporating higher-curvature gravitational dynamics into NR string heterotic backgrounds.
@@ 1 Introduction
Historically, these corrections have been explored via two main approaches. One is the amplitude-based method, whose Lagrangian was first constructed by the Metsaev-Tseytlin action [2], in agreement with the string scattering amplitudes [3]-[4]. The other approach, pioneered by Bergshoeff and de Roo, uses the restrictions given by supersymmetric invariance to construct higher-derivative completions of heterotic supergravity [5]. In particular, for the bosonic sector of the theory, Bergshoeff and de Roo found an identification which reveals a profound structural analogy between the gauge and the gravitational sector, offering a unified treatment of the -corrections of the theory. This identification has not only clarified the structure of the four-derivative gravitational corrections but has also provided a systematic way to encode the Green-Schwarz anomaly cancellation mechanism in terms of a torsionful spin connection.
@@
In the context of the low energy limit of non-relativistic string theory [6]-[13], the Bergshoeff–de Roo identification has not yet been explored. This extension to the NR regime provides a systematic framework for organizing -corrections in a manner consistent with NR symmetries. Since the leading-order NS-NS sector of string theory remains finite under a well-defined NR expansion [14] (due to precise cancellations between divergent terms) it is natural to extend this construction to the four-derivative gravitational corrections. In this work we achieve this goal by considering a suitable generalization of the Bergshoeff-de Roo identification over the two-derivative terms in the gauge sector (Chern-Simons plus Yang-Mills) of the theory [15] 111See also [16] and the appendix B in [15], for an alternative formulation of the gauge sector. We will discuss the Bergshoeff-de Roo identification of this alternative formulation in the discussion section.. The strategy is to construct the four-derivative gravitational terms from the two-derivative gauge theory, turning gauge contributions into gravitational contributions. For a pedagogical introduction to this technique in the relativistic regime, see [17].
@@
The main goal of this paper is to construct the NR analogue of the Bergshoeff–de Roo identification and use it to derive the four-derivative corrections to the heterotic supergravity action in the NR limit. To this end, we begin by reviewing the relativistic Bergshoeff-de Roo mechanism (section 2), highlighting its reinterpretation of the gravitational -corrections as arising from a map between non-Abelian gauge fields and torsionful Lorentz connections. We then apply a systematic NR expansion to the bosonic part of the two-derivative heterotic action, including both the NS-NS and gauge sectors and construct the NR version of the Bergshoeff-de Roo identification (section 3). The key technical tool is the light-cone decomposition of the Newton-Cartan geometry, which allows us to isolate the degrees of freedom responsible for the non-trivial transformations of the Kalb-Ramond two-form under local Lorentz transformations. By implementing a field redefinition that isolates the longitudinal component in the Green-Schwarz mechanism, we match the structure of the non-Abelian gauge transformation to the form of the gravitational anomaly cancellation. This redefinition is inspired by the recent results in [18]. Since the transformation is ambiguous in this limit, instead of eliminating the whole transformation one can remove only the longitudinal component, which is the key step implemented to perform the Bergshoeff-de Roo identification. This leads us to identify the NR gauge field with an spin connection with torsion, and to express the Chern-Simons-like contributions to the three-form field strength in terms of the gravitational degrees of freedom.
@@
We then apply this identification to the finite action obtained after taking the NR limit of heterotic supergravity. The resulting four-derivative action, detailed in Section 4 and summarized in Section 1.1, contains finite terms depending on the curvature and on the intrinsic torsion . By construction the full action is invariant under , boosts transformations and diffeomorphism invariance, the first one provided by the Bergshoeff de Roo identification. Particularly, the new four-derivative terms are not invariant, but they compensate the transformation of the two-derivative action with respect to the Green-Schwarz mechanism, which can be trivialized with field redefinitions.
@@
In section 5 we discuss two possible immediate continuations for this paper: on the one hand, the construction of the Metsaev and Tseytlin action [2] in the NR limit. This formulation is promising because it could provide the form of the four-derivative bosonic corrections, which are not guided by a Green-Schwarz mechanism. On the other hand, we also discuss the construction of the four-derivative corrections associated with the formulation [16]. In section 5.2 we obtain the form of the Green-Schwarz mechanism using the Bergshoeff-de Roo identification for a particular case of this construction, and it would be interesting to extend this result to obtain the corrections for this prescription. Finally, in section 6, we conclude the work and discuss some potential long-term lines of research.
1.1 Main result
The main result of this paper is the explicit construction of the bosonic sector of the heterotic supergravity Lagrangian up to four-derivative terms, in the NR limit of the theory (),
where
and
with a gauge vector, and the gauge and invariant 3-form given by
The covariant derivatives contain both a gauge connection and a connection with torsion defined as
Finally, the transformations of the fundamental fields are
where is a parameter of infinitesimal diffeomorphisms, is a parameter of non-Abelian gauge transformations, is a boost parameter, and is a parameter of transformations. The action of the heterotic supergravity up to four-derivative terms is invariant under all these symmetries. We recall that the final contribution to the transformation can be eliminated by using field-redefinitions, as shown in [18]. Imposing this redefinition would simplify the transformations and makes them compatible with the transformations of the bosonic supergravity in the Metsaev and Tseytlin approach [2] (see section 5.1 for a discussion on this topic). The invariance of the action with respect to and dilation transformations [14], which are present in the leading order, are not preserved by the identification process. An alternative identification to preserve these symmetries is given in the section 5.2, by applying the identification in the a particular case of the formulation given in [16], and the price to pay is that the Green-Schwarz mechanism cannot be trivialized. While this is not a general feature of the prescription given in [16], further study is required to compatibilize the Bergshoeff-de Roo identification in this setup and in section 5.2 we just limitate the study to the case where the expansion of the vielbein does not contain gauge contributions, as discussed in the appendix B of [15].
@@ 3 The non-relativistic version of the Bergshoeff-de Roo identification
where are space-time indices and we have split the flat index where are the longitudinal directions and the transverse ones. The quantities in the previous expansion obey the following Newton-Cartan relations,
While, a priori, this expansion produces divergent tensorial quantities upon taking the NR limit (i.e. Riemann tensor and its traces, H strength, etc.), the NS-NS supergravity Lagrangian,
@@
with . In this case, the Lagrangian also remains finite after considering the limit [15] and it can be written as
@@
While the previous transformations take the usual structure of heterotic supergravity with acting as a connection, the expansion (3.11) breaks the and the dilation symmetry explicitly and, therefore, the four-derivative gravitational correction constructed by the Bergshoeff-de Roo identification will inherit this effect (for an alternative identification preserving these symmetries see section 5.2).
On the other hand, the gravitational and torsion full Green-Schwarz mechanism after taking the limit is given by 222This expression includes the torsion in the SO(8) Green-Schwarz mechanism and agrees with the result in [18], which only contains the torsionless part.,
@@ 4 The gravitational non-relativistic heterotic action up to four-derivatives
We impose the Bergshoeff-de Roo identifications in the gauge contributions of (3.17) and (3.19). Let us start by finding the form of after the identifications,
@@
Since the previous contribution already has two-derivatives, it is not necessary to compute since this contribution will correspond to a contribution which we will neglect. However, as we can see, the procedure is predicting a higher-order structure above the order, similarly to what happens in the relativistic case.
@@
while the four-derivative corrections related to the Riemann tensor are given by
where the covariant derivatives now contain the connection, providing covariant quantities under this symmetry, together with the gravitational Green-Schwarz mechanism.
@@ 5.1 Towards a Metsaev-Tseytlin formulation of non-relativistic heterotic string
The four-derivative corrections to the universal action principle of the different formulations of closed string theory were historically computed considering three- and four-point scattering amplitudes for the massless states [3] [4]. This method is based on the study of the different string interactions through the S-matrix, in order to construct an effective Lagrangian which would be able to reproduce that interactions. The effective action, originally computed by Metsaev and Tseytlin, takes the form
@@
While in the relativistic case both formulations are physically the same (see [20] for the full proof), the non-relativistic limit of the action (5.2) has not yet been obtained. The Riemann squared contribution contains divergences [21] that should be canceled by the Chern-Simons contributions as happens in the formulation constructed here using the Bergshoeff-de Roo identification. For example, our Lagrangian does not contain terms given by since in the two-derivative heterotic Lagrangian we do not have terms after considering the NR limit. This cancellation of divergences in the Metsaev and Tseytlin formulation in the NR limit could be related to field redefinitions equivalent to (5.6) in this limit or to higher-derivative extensions of the non-relativistic expansion of the fundamental fields. Finally, the construction of the Metsaev and Tseytlin action could allow us to obtain the form of the gravitational four-derivative corrections of the bosonic string, so it integrates a relevant future direction of research.
@@ 5.2 Alternative identification: Bergshoeff-Romano formalism
Part of the finite contributions to the gravitational action in the Bergshoeff-Romano formulation are given by the squared of the previous object, while other contributions should be constructed by considering the identification over the full finite action (see appendix B of [15]). The advantage of this formulation is that the identification process respects the and dilatation invariance of the finite gauge action, while the disadvantage, in the case , is that the gauge Green-Schwarz mechanism cannot be trivialized. Therefore, relaxing the and find the complete form of the gravitational corrections is a promising way to further develop the identification in this formalism.
@@ 6 Conclusions and outlook
Using this framework, we derived the finite four-derivative corrections to the NR heterotic action. The resulting action is manifestly invariant under ( transformation, local boost transformations and diffeomorphism invariance. While the leading two-derivative NSNS action enjoys scale invariance and , the gauging with breaks these symmetries and, consequently, they are not present in the four-derivative terms generated via the Bergshoeff-de Roo procedure. Nonetheless, this symmetry breaking may be avoided by performing non-covariant field redefinitions to connect, for example, the two-derivative formulation with the one in [16]. The exploration of interesting topic is left for future work.
@@
T-duality invariant rewriting: Finally, one interesting possibility is to uplift the present formulation in the framework of Double Field Theory [23]-[24], in a non-Riemannian formulation [25]-[29], with higher-derivative terms following the construction given in [30]. It would be very interesting to incorporate the technology of the generalized Bergshoeff-de Roo identification constructed in [31]-[33] in the NR limit of Double Field Theory given in [19].

### v2

@@ Abstract
We present the first explicit construction of the four-derivative gravitational corrections to heterotic supergravity in the non-relativistic (NR) limit. By extending the Bergshoeff–de Roo identification to NR backgrounds, we obtain the full finite four-derivative completion of the gravitational sector of the NR heterotic supergravity action. A key outcome is the emergence of a gravitational Green–Schwarz mechanism from the NR gauge transformation of the Kalb–Ramond field. This mechanism can be trivialized through suitable field redefinitions, in agreement with a previous analysis of this transformation. Our results establish a systematic framework for incorporating higher-curvature gravitational dynamics into NR string heterotic backgrounds.
@@ 1 Introduction
Historically, these corrections have been explored via two main approaches. One is the amplitude-based method, whose Lagrangian was constructed by Metsaev-Tseytlin [2], in agreement with the string scattering amplitudes [3]-[4]. The other approach, pioneered by Bergshoeff and de Roo, uses the restrictions imposed by supersymmetric invariance to construct higher-derivative completions of heterotic supergravity [5]. In particular, for the bosonic sector of the theory, Bergshoeff and de Roo found an identification which reveals a profound structural analogy between the gauge and the gravitational sector, offering a unified treatment of the -corrections of the theory. This identification has not only clarified the structure of the four-derivative gravitational corrections but has also provided a systematic way to encode the Green-Schwarz anomaly cancellation mechanism in terms of a torsionful spin connection.
@@
In the context of the low energy limit of non-relativistic string theory [6]-[13], the Bergshoeff–de Roo identification has not yet been explored. This extension to the NR regime provides a systematic framework for organizing -corrections in a manner consistent with NR symmetries. Since the leading-order NS-NS sector of string theory remains finite under a well-defined NR expansion [14] (due to precise cancellations between divergent terms in the Lagrangian),
where are spacetime indices and we have split the flat index where are the longitudinal directions and are the transverse ones. The quantities in the previous expansion obey the following Newton-Cartan relations,
It is natural to wonder about extensions of this construction to the four-derivative corrections. In this work we focus on the gravitational sector of this corrections, and construct the four-derivative Lagrangian by considering a suitable generalization of the Bergshoeff-de Roo identification. We impose this map over the two-derivative terms in the gauge sector (Chern-Simons plus Yang-Mills) of the theory [15] 111We also impose the Bergshoeff-de Roo identification on the pioneer heterotic formulation constructed [16]. Since in this formulation transforms under gauge transformations, we limit the identification to the particular case where this transformation is trivial.. The strategy is to construct the four-derivative gravitational terms from the two-derivative gauge theory, turning gauge contributions into gravitational contributions. For a pedagogical introduction to this technique in the relativistic regime, see [17].
@@
The main goal of this paper is to construct the NR analogue of the Bergshoeff–de Roo identification and use it to derive the four-derivative corrections to the heterotic supergravity action in the NR limit. To this end, we begin by reviewing the relativistic Bergshoeff-de Roo mechanism (section 2), highlighting its reinterpretation of the gravitational -corrections as arising from a map between non-Abelian gauge fields and torsionful Lorentz connections. We then apply a systematic NR expansion to the bosonic part of the two-derivative heterotic action, including both the NS-NS and gauge sectors,
where , and we construct the NR version of the Bergshoeff-de Roo identification (section 3). The key technical tool is the light-cone decomposition of the Newton-Cartan geometry, which allows us to isolate the degrees of freedom responsible for the non-trivial transformations of the Kalb-Ramond two-form under local Lorentz transformations. By implementing a field redefinition that isolates the longitudinal component in the Green-Schwarz mechanism, we match the structure of the non-Abelian gauge transformation to the form of the gravitational anomaly cancellation. This redefinition is inspired by the recent results in [18]. Since the transformation is ambiguous in this limit, instead of eliminating the whole transformation one can remove only the longitudinal component, which is the key step implemented to perform the Bergshoeff-de Roo identification. This leads us to identify the NR gauge field with an spin connection with torsion, and to express the Chern-Simons-like contributions to the three-form field strength in terms of the gravitational degrees of freedom.
@@
We then apply this identification to the finite action obtained after taking the NR limit of heterotic supergravity. The resulting four-derivative action, detailed in Section 4, contains finite terms depending on the curvature and on the intrinsic torsion . By construction the full action is invariant under , boost transformations and diffeomorphism invariance, the first one provided by the Bergshoeff-de Roo identification. In particular, the new four-derivative terms are not invariant, but they compensate the transformation of the two-derivative action with respect to the Green-Schwarz mechanism, which can be trivialized with field redefinitions.
@@
In section 5 we discuss two possible immediate continuations for this paper: on the one hand, the construction of the Metsaev and Tseytlin action [2] in the NR limit. This formulation is promising because it could provide the form of the four-derivative bosonic corrections, which are not guided by a Green-Schwarz mechanism. On the other hand, we also discuss the construction of the four-derivative corrections associated with the formulation [16]. In section 5.2 we obtain the form of the Green-Schwarz mechanism using the Bergshoeff-de Roo identification for a particular case of this construction, and we discuss the possibility of extending this approach to obtain the gravitational four-derivative corrections for this prescription. Finally, in section 6, we conclude the work and discuss some potential long-term lines of research.
@@ 3 The non-relativistic version of the Bergshoeff-de Roo identification
While a priori this expansion produces divergent tensorial quantities upon taking the NR limit (i.e. Riemann tensor and its traces, H strength, etc.), the NS-NS supergravity Lagrangian,
@@
with . In this case, the Lagrangian also remains finite after taking the limit [15] and it can be written as
@@
On the other hand, the gravitational and torsionful Green-Schwarz mechanism after taking the limit is given by 222This expression includes the torsion in the SO(8) Green-Schwarz mechanism and agrees with the result in [18], which only contains the torsionless part.,
@@ 4 The gravitational non-relativistic heterotic action up to four-derivatives
We impose the Bergshoeff-de Roo identifications in the gauge contributions of (3.15) and (3.17). The expression for after the identifications is given by,
@@
Since the previous contribution already has two derivatives, it is not necessary to compute , this contribution will correspond to a contribution. However, as we can see, the procedure is predicting a higher-order structure above the order, similarly to what happens in the relativistic case.
@@
while the four-derivative corrections associated to the Riemann tensor are given by
where the covariant derivatives now contain the connection, providing covariant quantities under this symmetry, together with the gravitational Green-Schwarz mechanism. Finally, the transformations of the fundamental fields are
where is a parameter of infinitesimal diffeomorphisms, is a parameter and is a parameter of transformations. The action of heterotic supergravity up to four-derivative terms is invariant under all these symmetries plus boost transformations, which we have not explicitly included in the previous list since they would need higher-derivative contributions in the transformations. We recall that the final contribution to the transformation can be eliminated by using field-redefinitions, as shown in [18]. Imposing this redefinition would simplify the transformations and makes them compatible with the transformations of the bosonic supergravity in the Metsaev and Tseytlin approach [2] (see section 5.1 for a discussion on this topic). An alternative identification is given in the section 5.2, by applying the identification in the a particular case of the formulation given in [16], paying the price that the Green-Schwarz mechanism cannot be trivialized. While this is not a general feature of the prescription given in [16], further study is required to compatibilize the Bergshoeff-de Roo identification in this setup and in section 5.2 we just limit the study to the case where the expansion of the vielbein does not contain gauge contributions.
@@ 5.1 Towards a Metsaev-Tseytlin formulation of non-relativistic heterotic string
The four-derivative corrections to the universal action principle of the different formulations of closed string theory were historically computed considering three- and four-point scattering amplitudes for the massless states [3] [4]. This method is based on the study of the different string interactions through the S-matrix, in order to construct an effective Lagrangian which would be able to reproduce those interactions. The effective action, originally computed by Metsaev and Tseytlin, takes the form
@@
While in the relativistic case both formulations are physically the same (see [20] for the full proof), the non-relativistic limit of the action (5.2) has not yet been obtained. The Riemann squared contribution contains divergences [21] that should be canceled by the Chern-Simons contributions as happens in the formulation constructed here using the Bergshoeff-de Roo identification. For example, our Lagrangian does not contain terms given by since in the two-derivative heterotic Lagrangian we do not have terms after considering the NR limit. This cancellation of divergences in the Metsaev and Tseytlin formulation in the NR limit could be related to field redefinitions equivalent to (5.6) in this limit or to higher-derivative extensions of the non-relativistic expansion of the fundamental fields. Finally, the construction of the Metsaev and Tseytlin action could allow us to obtain the form of the gravitational four-derivative corrections for the bosonic string, so it integrates a relevant future direction of research.
@@ 5.2 Alternative identification: Bergshoeff-Romano formalism
Part of the finite contributions to the gravitational action in the Bergshoeff-Romano formulation are given by the squared of the previous object, while other contributions should be constructed by considering the identification over the full finite action (see appendix B of [15]). The disadvantage of this alternative expansion in the case , is that the gauge Green-Schwarz mechanism cannot be trivialized. Therefore, relaxing the and find the complete form of the gravitational corrections is a promising way to further develop the identification in this formalism.
@@ 6 Conclusions and outlook
Using this framework, we derived the finite four-derivative corrections to the NR heterotic action. The resulting action is manifestly invariant under transformation, transformations and diffeomorphism invariance. While the leading two-derivative NSNS action enjoys boost invariance, the four-derivative terms generated via the Bergshoeff-de Roo procedure requires higher-derivative corrections to these transformations. While the boost symmetry is guaranteed due to the Double Field Theory (DFT) rewriting of the gauge action, the exploration of the final form of these transformation is left for future work.
@@
Inclusion of boost transformations: The construction of the higher-derivative boost transformations can be systematically done from Heterotic DFT. Since the generalized metric is an element of , with the dimension of the heterotic gauge group, the boost transformations of the gauge degrees of freedom can be read from the transformations of the components of this metric. While the authors in [19] have not explicitly include these transformations in their analysis, it would be a straightforward computation to obtain those transformations and apply the non-relativistic version of the Bergshoeff-de Roo map on them. Since this symmetry is guaranteed due to the generalized metric formalism, it would be interesting to explore the explicit form of the transformations.
Inclusion of Supersymmetry: Another possibility is to incorporate the fermionic sector of the theory, following the lines of the original formulation of Bergshoeff and de Roo [5], in the NR limit. Initial steps in this direction were given in [22] by considering the leading order in fermions. While the supersymmetry transformations diverge in this prescription, starting from DFT could bring new possibilities to explore these transformations using a more general set of fields.
T-duality invariant rewriting: Finally, one interesting possibility is to uplift the present formulation in the framework of DFT [23]-[24], in a non-Riemannian formulation [25]-[29], with higher-derivative terms following the construction given in [30]. It would be very interesting to incorporate the technology of the generalized Bergshoeff-de Roo identification constructed in [31]-[34] 333See also [35] for a recent new approach to this topic. in the NR limit of DFT given in [19].